# SEOE: A Scalable and Reliable Semantic Evaluation Framework for Open Domain Event Detection

### v1

@@ Abstract
Automatic evaluation for Open Domain Event Detection (ODED) is a highly challenging task, because ODED is characterized by a vast diversity of un-constrained output labels from various domains. Nearly all existing evaluation methods for ODED usually first construct evaluation benchmarks with limited labels and domain coverage, and then evaluate ODED methods using metrics based on token-level label matching rules. However, this kind of evaluation framework faces two issues: (1) The limited evaluation benchmarks lack representatives of the real world, making it difficult to accurately reflect the performance of various ODED methods in real-world scenarios; (2) Evaluation metrics based on token-level matching rules fail to capture semantic similarity between predictions and golden labels. To address these two problems above, we propose a scalable and reliable Semantic-level Evaluation framework for Open domain Event detection (SEOE) by constructing a more representative evaluation benchmark and introducing a semantic evaluation metric. Specifically, our proposed framework first constructs a scalable evaluation benchmark that currently includes 564 event types covering 7 major domains, with a cost-effective supplementary annotation strategy to ensure the benchmark’s representativeness. The strategy also allows for the supplement of new event types and domains in the future. Then, the proposed SEOE leverages large language models (LLMs) as automatic evaluation agents to compute a semantic F1-score, incorporating fine-grained definitions of semantically similar labels to enhance the reliability of the evaluation. Extensive experiments validate the representatives of the benchmark and the reliability of the semantic evaluation metric. Existing ODED methods are thoroughly evaluated, and the error patterns of predictions are analyzed, revealing several insightful findings. Our benchmark and evaluation toolkits are publicly available.
@@ 1 Introduction
Figure 1: Comparison between previous evaluation frameworks and our proposed SEOE.
Open domain event detection (ODED) is a crucial task in open domain information extraction. Its primary objective is to extract structured information, namely events, from large-scale textual data across various domains without constrained ontologies. This structured information can then be used for further research tasks, such as information retrieval Zhang et al. (2024); Joshi et al. (2023), knowledge base construction Zhao et al. (2024), and large language models Cai et al. (2024); Chen et al. (2024); Wang et al. (2023), where understanding the relationships between events is essential. Recently, the importance of ODED has increased with the rapid expansion of the web, which continuously generates vast amounts of unstructured information across various domains, such as clinical area Ma et al. (2023a), suicide understanding Guzman-Nateras et al. (2022), and scientific publications Lu et al. (2024b).
Evaluating ODED is a highly challenging task due to the vast diversity of event types in the open domain Araki and Mitamura (2018). When making predictions, ODED models are often required to extract events with unseen event types during the training and validation stage Cai et al. (2024). To evaluate these models, to the best of our knowledge, nearly all existing ODED evaluation frameworks simulate the open setting by evaluating models using a predefined ontology in a zero-shot manner Cai et al. (2024); Yue et al. (2023). Specifically, as illustrated in Figure 1, these frameworks typically first construct or utilize evaluation benchmarks with a limited set of types and domain coverage that are not seen during the training phase. During evaluation, ODED models are required to recognize events constrained by predefined ontologies of these benchmarks. These evaluation frameworks then apply token-level matching rules to evaluate models’ predictions, such as strict match Zhang et al. (2022a), partial match Araki and Mitamura (2018), and head noun match Li et al. (2021).
@@
To address these two problems above, we propose a scalable and reliable Semantic Evaluation framework for Open domain Event detection (SEOE). This framework consists of a more representative evaluation benchmark and a semantic evaluation metric. As for the benchmark, our proposed framework SEOE constructs a scalable evaluation benchmark that currently includes 564 event types covering 7 major domains. We first integrate ontologies of mainstream event extraction datasets into a larger ontology. Then, we propose a cost-effective supplementary annotation strategy to ensure that each data instance is examined by the integrated ontology, making the benchmark both representative of real-world scenarios and easily scalable. As for the evaluation metric, SEOE leverages large language models (LLMs) as automatic evaluation agents to compute a semantic F1-score. To ensure reliable evaluation, SEOE generates fine-grained definitions of event types and groups semantically similar labels together to assist in the evaluation process.
@@ 2.2 LLMs for Automatic Benchmark Construction
With the emergence of advanced LLMs, exemplified by GPT-4, automated data annotation and synthesis have garnered significant attention in the NLP community Tan et al. (2024). Many studies on LLM-based annotation focus on automating the labeling process with the help of LLMs Yadav et al. (2024); Tseng et al. (2024), including single LLM annotation Chen et al. (2024), joint annotation by multiple LLMs Martorana et al. (2024), and hybrid annotation by humans and LLMs Li et al. (2023). In addition, the evaluation of annotations is also a crucial part of automated annotation, with mainstream studies including general evaluation with human comparisons Honovich et al. (2023), task-specific evaluations Chen et al. (2023b), and using LLMs as evaluators known as LLM-as-a-judge Wu et al. (2024); Zheng et al. (2023). In this paper, we propose a cost-effective method to instruct GPT-4o to construct a representative ODED benchmark.
@@ 2.3 Evaluation for Event Detection
ED is typically evaluated using token-level matching rules Li et al. (2021); Lin et al. (2020), which assess whether the model-generated events and their labels are token-level identical or similar to the human-annotated ground truth. However, these rules exhibit weak correlation with human judgments due to the lack of the semantic evaluation Lu et al. (2024a). Recently, there is an emerging trend in the natural language generation research community to leverage LLMs for evaluating generations in a semantic manner Lan et al. (2024a, 2020). Notably, (Lu et al., 2024a) propose a LLM-based automatic evaluation framework for event extraction, named RAEE. However, RAEE is designed for closed-domain scenarios and faces the challenge of evaluating a wide range of event types with varying definitions in open-domain scenarios. In this paper, we propose a LLM-based automatic evaluation framework that includes a scalable benchmark and a reliable evaluation metric to semantically evaluate ODED models.
@@ 4.1 Benchmark Construction
Figure 3: Distribution of SEOE benchmark
@@ 4.2 Benchmark Reliability Enhanced by Nucleus Sampling
Table 1: The results of different annotation strategies on the number of supplementary triggers (Trg. Count) and annotation accuracy (Correct Trg.).
To validate the above strategy’s effectiveness and the benchmark’s reliability, we sample 200 data instances and conduct a quality check by 3 humans on the supplementary annotation data, as shown in Table 1. Results show that (1) With a fixed $p$, increasing annotation rounds improves both the number and accuracy of supplementary triggers. This demonstrates that multiple annotations and nucleus sampling strategy are highly effective in enhancing the benchmark’s reliability. (2) When annotation rounds are fixed, an increase in $p$ leads to a rise in the total number of supplementary annotations, but the annotation accuracy decreases. This suggests that adjusting $p$ represents a trade-off between the diversity and accuracy of the supplementary annotations. We will release three versions of the evaluation benchmark, based on 10 rounds of repeated annotations and using $p=0.3$, $p=0.5$, and $p=0.7$, allowing future research to choose an evaluation that favors accuracy ($p=0.3$) or diversity ($p=0.7$). In subsequent experiments, we will consistently use the balanced version ($p=0.5$) for evaluation.
@@ 4.3 Benchmark Statistics and Scalability
Since ODED aims to detect all possible events from any potential domain, it is essential to ensure the scalability of the ODED evaluation benchmark. In SEOE, both data and event types are scalable. Specifically, when new data is added to the benchmark, step 3 and 4 are repeated to achieve new annotations. When new event types are added, step 2 generates their fine-grained definitions. Then, step 3 and 4 perform new supplementary annotations for previous data. Since step 3 employs a text similarity model for filtering, the additional cost incurred from adding new data and event types is approximately linear with the number of data instances and event types. This ensures the benchmark’s cost-effective scalability and enhances the model’s representativeness of real-world scenarios. To support this conclusion, a formal analysis can be found in Appendix D.2.
@@ 5.3 LLMs’ Evaluation Correlation with Human Judgments
To prove the effectiveness of our proposed SEOE framework and several modules, we conduct a meta-evaluation to analyze the correlation between SEOE and human judgments. Following previous work’s setting Lu et al. (2024a), we invite 3 professional event annotators to evaluate 200 sampled pieces of data. These annotators are given the same prompts as those given to LLMs. As for the evaluation agents, we select the top-3 LLMs that demonstrate superior performance in text evaluation [^2] https://openai.com/index/hello-gpt-4o/. As for the evaluation metrics, we compute three widely-used inter-annotator agreement (IAA) metrics among humans and LLMs: percent agreement Zheng et al. (2023); Liu et al. (2021b), Spearman correlation Lan et al. (2024b), and Cohen’s Kappa Thakur et al. (2024). To minimize the impact of randomness from LLMs, we repeat the experiment three times and report the average and standard deviation of the results, as shown in Table 2.
@@ 7 Conclusion and Future Works
Unlike evaluations based on token-level matching rules, our evaluation framework utilizes advanced LLMs to assess models’ ODED predictions, which introduces higher costs. In the case of the full benchmark, which contains 4,423 data instances, the cost of evaluating a model is approximately \$59.8. However, it is possible to accurately evaluate models using only a subset of the benchmark, as the average frequency of each event type in our benchmark is approximately 50.8, which may exceed the requirements for model evaluation. To validate this hypothesis, we identify subsets of the benchmark with sizes of 1/10, 1/4, and 1/2 through random sampling. The average differences in evaluation metrics between these subsets and the full benchmark used in the main experiments are 1.06%, 0.57%, and 0.38%, respectively. This indicates that using an appropriately chosen subset can significantly reduce evaluation costs while maintaining statistical representativeness. These subsets of the benchmark will also be publicly available.

### v2

@@ Abstract
Automatic evaluation for Open Domain Event Detection (ODED) is a highly challenging task, because ODED is characterized by a vast diversity of un-constrained output labels from various domains. Nearly all existing evaluation methods for ODED usually first construct evaluation benchmarks with limited labels and domain coverage, and then evaluate ODED methods using metrics based on token-level label matching rules. However, this kind of evaluation framework faces two issues: (1) The limited evaluation benchmarks lack representatives of the real world, making it difficult to accurately reflect the performance of various ODED methods in real-world scenarios; (2) Evaluation metrics based on token-level matching rules fail to capture semantic similarity between predictions and golden labels. To address these two problems above, we propose a scalable and reliable Semantic-level Evaluation framework for Open domain Event detection (SEOE) by constructing a more representative evaluation benchmark and introducing a semantic evaluation metric. Specifically, our proposed framework first constructs a scalable evaluation benchmark that currently includes 564 event types covering 7 major domains, with a cost-effective supplementary annotation strategy to ensure the benchmark’s representativeness. The strategy also allows for the supplement of new event types and domains in the future. Then, the proposed SEOE leverages large language models as automatic evaluation agents to compute a semantic F1-score, incorporating fine-grained definitions of semantically similar labels to enhance the reliability of the evaluation. Extensive experiments validate the representatives of the benchmark and the reliability of the semantic evaluation metric. Existing ODED methods are thoroughly evaluated, and the error patterns of predictions are analyzed, revealing several insightful findings. Our benchmark and evaluation toolkits are publicly available.
@@ 1 Introduction
Figure 1: Comparison between previous ODED evaluation frameworks and our proposed SEOE.
Open domain event detection (ODED) is a crucial task in open domain information extraction. Its primary objective is to extract structured information, namely events, from large-scale textual data across various domains without constrained ontologies. This structured information can then be used for further research tasks, such as information retrieval Zhang et al. (2024); Joshi et al. (2023), knowledge base construction Zhao et al. (2024), and large language models Cai et al. (2024); Chen et al. (2024); Wang et al. (2023), where understanding the relationships between events is essential. Recently, the importance of ODED has increased with the rapid expansion of the web, which continuously generates vast amounts of unstructured information across various domains, such as the clinical area Ma et al. (2023a), suicide understanding Guzman-Nateras et al. (2022), and scientific publications Lu et al. (2024b).
Evaluating ODED is a highly challenging task due to the vast diversity of event types in the open domain Araki and Mitamura (2018). When making predictions, ODED models are often required to extract events with unseen event types during the training and validation stage Cai et al. (2024). To evaluate these models, to the best of our knowledge, nearly all existing ODED evaluation frameworks simulate the open setting by evaluating models using a predefined ontology in a zero-shot manner Cai et al. (2024); Yue et al. (2023). Specifically, as illustrated in Figure 1, these frameworks typically first construct or utilize evaluation benchmarks with a limited set of types and domain coverage that are unseen during the training stage. During the evaluation process, ODED models are required to recognize events constrained by predefined ontologies of these benchmarks. These evaluation frameworks then apply token-level matching rules to evaluate models’ predictions, such as strict match Zhang et al. (2022a), partial match Araki and Mitamura (2018), and head noun match Li et al. (2021).
@@
To address these two problems above, we propose a scalable and reliable Semantic Evaluation framework for Open domain Event detection (SEOE). This framework consists of a more representative evaluation benchmark and a semantic evaluation metric. As for the benchmark, our proposed framework SEOE constructs a scalable evaluation benchmark that currently includes 564 event types covering 7 major domains. We first integrate the ontologies of mainstream event extraction datasets into a larger ontology. Then, we propose a cost-effective supplementary annotation strategy to ensure that each data instance is examined by the integrated ontology, making the benchmark both representative of real-world scenarios and easily scalable. As for the evaluation metric, SEOE leverages large language models (LLMs) as automatic evaluation agents to compute a semantic F1-score. To ensure reliable evaluation, SEOE generates fine-grained definitions of event types and groups semantically similar labels together to assist in the evaluation process.
@@ 2.2 LLMs for Automatic Benchmark Construction
With the emergence of advanced LLMs, exemplified by GPT-4, automated data annotation and synthesis have garnered significant attention in the NLP community Tan et al. (2024). Many studies on LLM-based annotation focus on automating the labeling process with the help of LLMs Yadav et al. (2024); Tseng et al. (2024), including single LLM annotation Chen et al. (2024); Xu et al. (2025), joint annotation by multiple LLMs Martorana et al. (2024), and hybrid annotation by humans and LLMs Li et al. (2023). In addition, the evaluation of annotations is also a crucial part of automated annotation, with mainstream studies including general evaluation with human comparisons Honovich et al. (2023), task-specific evaluations Chen et al. (2023b), and using LLMs as evaluators known as LLM-as-a-judge Wu et al. (2024); Zheng et al. (2023). In this paper, we propose a cost-effective method to instruct GPT-4o to construct a representative ODED benchmark.
@@ 2.3 Evaluation for Event Detection
ED is typically evaluated using token-level matching rules Li et al. (2021); Lin et al. (2020), which assess whether the model-generated events and their labels are token-level identical or similar to the human-annotated ground truth. However, these rules exhibit weak correlation with human judgments due to the lack of the semantic evaluation Lu et al. (2024a). Recently, there is an emerging trend in the natural language generation research community to leverage LLMs for evaluating generations in a semantic manner Li et al. (2025); Lan et al. (2024a, 2020). Notably, (Lu et al., 2024a) propose a LLM-based automatic evaluation framework for event extraction, named RAEE. However, RAEE is designed for closed-domain scenarios and faces the challenge of evaluating a wide range of event types with varying definitions in open-domain scenarios. In this paper, we propose a LLM-based automatic evaluation framework that includes a scalable benchmark and a reliable evaluation metric to semantically evaluate ODED models.
@@ 4.1 Benchmark Construction
Figure 3: Distribution of SEOE benchmark.
@@ 4.2 Benchmark Reliability Enhanced by Nucleus Sampling
Table 1: The results of different annotation strategies on the number of supplementary triggers (Trg. Count) and annotation accuracy (Correct Trg.). There is a trade-off between the diversity and accuracy of the supplementary annotations.
To validate the above strategy’s effectiveness and the benchmark’s reliability, we sample 200 documents (2,277 events in total) and conduct a quality check by 3 humans on the supplementary annotation data, as shown in Table 1. Results show that (1) With a fixed $p$, increasing annotation rounds improves both the number and accuracy of supplementary triggers. This demonstrates that multiple annotations and nucleus sampling strategy are highly effective in enhancing the benchmark’s reliability. (2) When the number of annotation rounds is fixed, an increase in $p$ leads to a rise in the total number of supplementary annotations, but the annotation accuracy decreases. This suggests that adjusting $p$ represents a trade-off between the diversity and accuracy of the supplementary annotations. A more detailed analysis can be found in Appendix C. We will release three versions of the evaluation benchmark, based on 10 rounds of repeated annotations and using $p=0.3$, $p=0.5$, and $p=0.7$, allowing future research to choose an evaluation that favors accuracy ($p=0.3$) or diversity ($p=0.7$). In subsequent experiments, we will consistently use the balanced version ($p=0.5$) for evaluation.
@@ 4.3 Benchmark Statistics and Scalability
Since ODED aims to detect all possible events from any potential domain, it is essential to ensure the scalability of the ODED evaluation benchmark. In SEOE, both data (i.e., event-containing documents) and event types are scalable. Specifically, when new data is added to the benchmark, step 3 and 4 are repeated to achieve new annotations. When new event types are added, step 2 generates their fine-grained definitions. Then, step 3 and 4 perform new supplementary annotations for previous data. Since step 3 employs a text similarity model for filtering, the additional cost incurred from adding new data and event types is approximately linear with the number of data instances and event types. This ensures the benchmark’s cost-effective scalability and enhances the model’s representativeness of real-world scenarios. To support this conclusion, a formulaic analysis can be found in Appendix D.2.
@@ 5.3 LLMs’ Evaluation Correlation with Human Judgments
To prove the effectiveness of our proposed SEOE framework and several modules, we conduct a meta-evaluation to analyze the correlation between SEOE and human judgments. Specifically, we invite 3 professional event annotators to evaluate 200 event-containing documents (791 prediction-gold event pairs in total). These annotators are given the same prompts as those given to LLMs. As for the evaluation agents, we select the top-3 LLMs that demonstrate superior performance in text evaluation [^2] https://openai.com/index/hello-gpt-4o/. As for the evaluation metrics, we compute three widely-used inter-annotator agreement (IAA) metrics among humans and LLMs: percent agreement Zheng et al. (2023); Liu et al. (2021b), Spearman correlation Lan et al. (2024b), and Cohen’s Kappa Thakur et al. (2024). To minimize the impact of randomness from LLMs, we repeat the experiment three times and report the average and standard deviation of the results, as shown in Table 2.
@@ 7 Conclusion and Future Works
Unlike evaluations based on token-level matching rules, our evaluation framework utilizes advanced LLMs to assess models’ ODED predictions, which introduces higher costs. In the case of the full benchmark, which contains 4,423 documents, the cost of evaluating a model is approximately \$59.8. However, it is possible to accurately evaluate models using only a subset of the benchmark, as the average frequency of each event type in our benchmark is approximately 50.8, which may exceed the requirements for model evaluation. To validate this hypothesis, we identify subsets of the benchmark with sizes of 1/10, 1/4, and 1/2 through random sampling. The average differences in evaluation metrics between these subsets and the full benchmark used in the main experiments are 1.06%, 0.57%, and 0.38%, respectively. This indicates that using an appropriately chosen subset can significantly reduce evaluation costs while maintaining statistical representativeness. These subsets of the benchmark will also be publicly available.